RC Meeting and Run Plan, Tue, January 29th 2020

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are at 4:15 PM, 2nd floor Counting House Meeting Room

  • Bluejeans remote connection information:
 To join the Meeting:
 To join via Room System:
 Video Conferencing System: bjn.vc -or-
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links

General updates

  • We are in the second half of the longitudinal polarization run (we did and NMR measurement ~12:54 - https://logbooks.jlab.org/entry/3773442). We will do another NMR measurement at 6 pm and then go to transverse polarization running. After 5 h of running in the transverse polarization configuration we will do another NMR and switch to longitudinal running. If the NMR measurement done before the transverse polarization running was 47 % or higher we will stay in the next longitudinal configuration running for 12 h.

Recent polarization measurements

 Long   51.5 %  12:54PM Wed 1/29 Mid way through the long polarization running period
 Trans  55%     10AM Tue 1/28   Rotated back to transverse because accelerator is down
 Long   56%     9AM Tue 1/28   After taking longitudinal data 
 Trans  61%     7AM Tue 1/28   After taking transverse data
 Long   51%    11PM Mon 1/27   After taking longitudinal data
 Trans  63%     3PM Mon 1/27   After taking transverse data
 Long   49%     8AM Mon 1/27   After taking longitudinal data
 Trans  57%     2AM Sun 1/26   After taking transverse data

Current Run Plan

  • Always take 120Hz of EDTM

Wednesday swing – Thursday day

!!                                                     !!
!!  ATTN: SHIFT LEADERS                                !!
!!     Please follow the shift summary template on     !!
!!     the laminated SL instructions.                  !!
!!                                                     !!
!!  ATTN: TARGET OPERATORS                             !!
!!  All NMR measurement HCLOG entries should include:  !!
!!     1) A title that clearly lists polarization      !!
!!     2) Screenshots of both NMR_FieldSweep and       !!
!!        the polarization fit                         !!
!!                                                     !!
!!     For an example see the following entry:         !!
!!     https://logbooks.jlab.org/entry/3768696       !!
!!                                                     !!


Moller & spin dance

  • Spin dance will be part of beam studies. At the scheduling meeting today it was proposed Wednesday next week as spin dance day.

Hall Access

  • None planned right now


  • Replace target cell in ~1 week

Special Requests

  • Keep target expert schedule filled out on whiteboard

Notices in Effect for Shift Crew

  • Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
  • Magnet policies
    • Spectrometer GUI (go_magnets) should always be visible on Shift Leader computer, hcdesk3
    • HMS Q1 must be ramped down in 50A steps in all cases
    • Do not ramp HMS Q2 to zero, only go to 10A and then switch off power.
    • If SHMS dipole trips again, take a screenshot of the PSU window before resetting the interlock and post it to hclog.
  • Inform the RC of any access that the Target Team requires. The RC will coordinate with RadCon to make everything run smoothly.
  • The operators have been somewhat careful with our beam orbit locks. Please make sure that the orbit locks are enabled and the beam does not drift from its nominal position.
  • Ramp rate lock instructions are here and added to checklist[1]
  • Temporary Laser instructions can be found here [2]
  • Do not change momentum in SHMS: Q3 right current lead valve needs to be fixed

Planned Accelerator Configuration Changes